How many bijections (bijective functions) are there from A to A with the

property that no element of A is mapped to itself? (A is a set with 5 elements)


1
Expert's answer
2022-05-10T13:27:43-0400

There are "5!" different bijections from "A" to "A". Namely the first element can be mapped into "5" different elements. The second one into "4" elements and so on. Using a multiplication principle of combinatorics, we receive: "5!=120" different bijections. We point out that a bijection that has "4" elements mapped to itself and "5" elements mapped to itself is the same. There is "1" bijection of this type. Consider bijections that have "3" elements mapped to itself, but not "4" and "5" . There are "C_5^3=\frac{5!}{3!2!}=10" maps of this type. Consider maps that have only "2" elements mapped to itself. There are "2C_5^2=2\frac{5!}{3!2!}=20" maps of this type. Consider maps that have "1" element mapped to itself. As an example, we fix "1". There are "4!" different bijections that have "1" mapped to itself. They contain "1" bijection that has "4" and "5" elements mapped to itself, "C_4^2=\frac{4!}{2!2!}=6" bijections that have "3" elements mapped to itself and "2C_4^1=8" bijections that have "2" elements mapped to itself. We receive: "4!-1-6-8=9" bijections. Thus, there are "9" bijections that have only "1" mapped to itself. We receive:"45" bijections that have only "1" element mapped to itself. Thus, we get: "1+10+20+45=76" different bijections that have at least "1" element mapped to itself. We receive: "120-76=44" bijections that have no elements mapped to itself.

Answer: there are "44" bijections that have no elements mapped to itself.